OpenStudy (anonymous):

graph the equation below plotting the y-intercept and a second point on the line. y=3x-5

OpenStudy (amistre64):

so what do you know about graphs and their equation?

OpenStudy (anonymous):

To find the y intercept, set X equal to 0 in the equation y= mx+b. So, first figure out what information you have.. looking at your equation, the y is unknown, m is equal to 3, x is unknown and b is equal to -5. y = 3x-5. The y intercept is found when you set x equal to 0, so replace x with a zero and rewrite your problem: y = (3)(0) - 5 Preform the calculation: 3 * 0 = 0 0 -5 = -5 y is equal to -5 and you know your x intercept is zero so you have your y intercept (0,-5)